Determine the mass of 4.99% AlF3 solution that contains 6.62 g of F-.
Answer #1

Moles of F- = mass of F / molar mass = 6.62 / 19 = 0.348 moles

Moles of AlF3 = moles of F- / 3 = 0.348 / 3 = 0.116 moles

Mass of AlF3 = moles * molar mass = 0.116 * 83.98 = 9.742 g

Mass of the solution = mass of AlF3 / mass % of salt * 100 = 9.742 * 100 / 4.99 = 195.23 g